OpenStudy (anonymous):

Log(base of)2(3x−1)=5 Not sure what to do after I get 31=3x Because I pretty sure you cant have a decimal

OpenStudy (photon336):

This is what you meant right? \[Log_{2}(3x-1) = 5\]

OpenStudy (anonymous):

Yes! I'm new and didn't know how to make equations like that

OpenStudy (photon336):

you see the equation button at the bottom of the typing text area? you can put in equations there

OpenStudy (anonymous):

Oh great. So do you know how I would go about solving this?

OpenStudy (anonymous):

Yeah, That is what I got. I don't know it doesn't seem correct

OpenStudy (anthonyym):

My bad I was looking at your answer. It should be 33 = 3x. You add to get rid of the -1

OpenStudy (photon336):

\[2^{5} = 3x-1 \] \[\frac{ (2^{5}+1) }{ 3 } = x | \frac{ 33 }{ 3 } = 11\]

OpenStudy (anonymous):

Oh my god...I don't know why I subtracted. I just noticed that

OpenStudy (anonymous):

Thanks alot :)

OpenStudy (photon336):

np :) welcome to OS @Bohdan3